OpenStudy (anonymous):

For the function f(x) = (8-2x)2 ,find f-1 . Determine whether f-1 is a function.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Astrophysics

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (campbell_st):

just start by swapping x and y so the original equation is \[y = (8 - 2x)^2\] do the swap \[x = (8 - 2y)^2\] now may y the subject of the equation

OpenStudy (anonymous):

So should i solve for x? @campbell_st

OpenStudy (campbell_st):

no you should solve for y... 1st step is to take the square root of both sides of the equation...

OpenStudy (anonymous):

y=4− x√ 2,4+ x√ 2

OpenStudy (campbell_st):

well the problem is that the original equation is a function, the inverse will be a relationship... and your solution seems ok...

OpenStudy (campbell_st):

the way you test the inverse is a function is to use the vertical line test after graphing the curve

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (anonymous):

It doesnt match any of my answers

OpenStudy (campbell_st):

so do you have \[f^{-1}(x) = \frac{8 - \sqrt{x}}{2}\]

OpenStudy (anonymous):

idk i cant see my answer A lol it's probably that one than

OpenStudy (anonymous):

All my other answers are a + not a - and I cant see answer A

OpenStudy (usukidoll):

take the one with the + ? Idk.. but @campbell_st is right. you switch the signs of x and y and solve for y

OpenStudy (anonymous):

Is the final answer correct? If so I'll just chose A because everything else is a + after the 8

OpenStudy (usukidoll):

?!!!! ok looks like I have to square root both sides or I won't go further lol

OpenStudy (anonymous):

haha I just want to make sure the answer is correct

OpenStudy (usukidoll):

\[x = (8 - 2y)^2\] [square both sides so we have y by itself] \[\sqrt{x}=8-2y\] ok so I'm just gonna stop at this point. I don't want to get busted :P but anyway we need y by itself so I ____________ 8 from both sides

OpenStudy (anonymous):

subtracr

OpenStudy (usukidoll):

yes \[\sqrt{x}-8=-2y\] so now what do we need to do to get y by itself?

OpenStudy (anonymous):

Add 2?

OpenStudy (usukidoll):

not quite. We can't use addition... try again

OpenStudy (anonymous):

divide?

OpenStudy (usukidoll):

yes we divide -2 on both sides.

OpenStudy (usukidoll):

\[\frac{\sqrt{x}-8}{-2}=y\] we can simplify further

OpenStudy (usukidoll):

for starters we can split this equation up.... \[\frac{\sqrt{x}-8}{-2}=y \rightarrow \frac{-\sqrt{x}}{2}+\frac{-8}{-2}\]

OpenStudy (anonymous):

So his answer was right than

OpenStudy (usukidoll):

so what's -8/-2 ?

OpenStudy (anonymous):

-4

OpenStudy (usukidoll):

a negative divided by a negative turns positive

OpenStudy (anonymous):

4

OpenStudy (usukidoll):

yeah.. and about his answer..I think he dropped the negative sign

OpenStudy (usukidoll):

\[\frac{-\sqrt{x}}{2}+4\]

OpenStudy (anonymous):

ok this is my final question tonight lol

OpenStudy (anonymous):

For the function f(x) = x2 - 12, find (f o f-1)(4)

OpenStudy (anonymous):

do we put 4 in all the x's?

OpenStudy (usukidoll):

hmmmm looks like we need to find f^{-1} first and plug that guy into f(x)

OpenStudy (usukidoll):

\[f(x) =x^2-12\] f(x) is also y \[y=x^2-12 \] now how do we find the inverse ?

OpenStudy (anonymous):

solve for y?

OpenStudy (usukidoll):

that's the second step what is the first step? what needs to be switched?

OpenStudy (anonymous):

the x and y

OpenStudy (usukidoll):

yes now we solve for y \[x=y^2-12\]

OpenStudy (usukidoll):

so how do we get y by itself? we need to __________ 12 to both sides

OpenStudy (anonymous):

add

OpenStudy (usukidoll):

\[x+12=y^2 \] good now we need to ___________ both sides we need a y by itself... it can't have an exponent

OpenStudy (anonymous):

divide?

OpenStudy (usukidoll):

no. umm.. what is the square root of y^2 ?

OpenStudy (anonymous):

nothing?

OpenStudy (usukidoll):

\[\sqrt{y^2} = ? \]

OpenStudy (usukidoll):

it can't be nothing.. otherwise we won't have y ._.

OpenStudy (usukidoll):

maybe exponential form will be easier \[\large y^\frac{2}{2}\] what's 2/2 ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

it would just turn into y

OpenStudy (usukidoll):

yes but keep in mind since we took the square root of y^2 we have to take the square root of the left hand side as well \[\sqrt{x+12}=y\]

OpenStudy (usukidoll):

eep \[f^{-1}(x) = \sqrt{x+12} \]

OpenStudy (usukidoll):

now we will place that inverse function into our \[f(x) = x^2-12 \] but something unusual is going to happen

OpenStudy (usukidoll):

oh wait it won't sorry we will still have something ok good whew

OpenStudy (anonymous):

haha ok

OpenStudy (usukidoll):

so f o f^{-1} means place f^{-1} inside the f(x) function.. yeah I was just looking at the number and not the variable xD

OpenStudy (anonymous):

Is the answer 6?

OpenStudy (usukidoll):

no try placing f inverse in f(x)

OpenStudy (anonymous):

4?

OpenStudy (usukidoll):

yeah

OpenStudy (anonymous):

sweetness

OpenStudy (anonymous):

im gonna need for you to be online whenever i am online lol

OpenStudy (usukidoll):

\[f^{-1}(x) = \sqrt{x+12} , f(x) = x^2-12 \] \[f(x) \cdot f^{-1}(x) \] \[f^{-1}(x) \cdot f(x) = (\sqrt{x+12})^2-12 \] \[f^{-1}(x) \cdot f(x) = x+12-12 \] \[f^{-1}(x) \cdot f(x) =x \] when x = 4 \[f^{-1}(x) \cdot f(x) =4 \]

OpenStudy (anonymous):

Thank you for all of your help, I got an A on my hw

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

oh that's awesome! :D